OpenStudy (anonymous):

If g(x) = 2x + 2, find g(a + h) - g(a).

OpenStudy (mayankdevnani):

\[\huge \color{red}{Welcome}~\color{green}{To}~\color{hotpink}{OpenStudy}\] @mikeoaster

OpenStudy (anonymous):

If g(x) = 2x + 2, find g(a + h) - g(a).

OpenStudy (mayankdevnani):

\[\huge g(x)=2x+2\] \[\huge g(a)=2a+2\] \[\huge g(a+h)=2(a+h)+2=2a+2h+2\]

OpenStudy (mayankdevnani):

then, \[\huge g(a+h)-g(a)=???\] can you solve it?? @mikeoaster

OpenStudy (anonymous):

already got it thanks

OpenStudy (mayankdevnani):

welcome:)

OpenStudy (mayankdevnani):

@mikeoaster you can give me medal also by clicking the BEST RESPONSE right side to my response

OpenStudy (mayankdevnani):

\[\huge clicking~best~response-->give~medal\]